OpenStudy (anonymous):

A ball is thrown vertically upward with initial velocity v. Find the maximum height H of the ball as a function of v. Then find the initial velocity v required to achieve a height of H.

OpenStudy (anonymous):

for the first part, what forces are acting upon the ball? only the initial velocity V or something else?

OpenStudy (anonymous):

So: vf^2 = vi^2 + 2ad At the top, final velocity is 0, so 0 = vi^2 + 2ad We know that the only acceleration is that of gravity, and that vi is just what you call "v" 0 = v^2 + 2(-9.8)(d) 19.6d = v^2 d = v^2 / 19.6 There you go!

OpenStudy (anonymous):

thats all the question says...its more a theoretically question... ive tried answering it a couples of ways (since our electronic submission allows for multiple entry)..

OpenStudy (anonymous):

i have tried v^2/19.6 and it says that is incorrect

OpenStudy (anonymous):

Is it asking you for a number or a formula?

OpenStudy (anonymous):

Height H achieved with given initial velocity v: Initial velocity v required to achieve given height H:

OpenStudy (anonymous):

im not sure...because it wont let me introduce another variable such as g ...says that is incorrect

OpenStudy (anonymous):

here is what I have, your height H is going to equal the initial velocity V - the force of gravity, which is mass times acceleration. H=V-ma

OpenStudy (anonymous):

it says that ma is not defined in this context unfortunately

OpenStudy (anonymous):

i am sorry, i dont quite know what its asking

OpenStudy (anonymous):

thats ok, thanks for trying to help though